Find the mole fractions in the liquid mixture and in the vapor phase

At 100 ?C the vapor pressures of hexane and octane are 1836 and 354 torr, respectively. A certain liquid mixture of these two compounds has a vapor pressure of 666torr at 100 ?C. Find the mole fractions in the liquid mixture and in the vapor phase. Assume an ideal solution.
Answer #1
1836 x + 354 ( 1-x ) = 666

x = 4/19

Hexane = 4/19
Octane = 15/19
answered by: Keishante